Consider the circuit shown below. Where R1 = 30 Ω , R2 = 40 Ω , and R3 = 90 Ω . When a battery with unknown voltage is hooked up between points A and B, the current through the resistor R1 is I1 = 0.050 A

R2 3

What is the voltage across R3? What is the current through the resistor R3?

What is the total current produced by the battery hooked between points A and B?

What is the voltage of the battery hooked to A and B?
Answer #1

Both R1 and R2 are in series so, current will be same for both resister

Current through R2 = 0.05 A

Voltage across R1 = 30*0.05 = 1.5 V

Voltage across R2 = 40*0.05 = 2 V

So voltage across R3 should be 1.5 + 2 = 3.5 V

Current through R3 = V/R = 3.5/90 = 0.039 A

Total current produce by battery = 0.05+0.039 = 0.089 A

Voltage of battery = 3.5 Volt